A right triangle has sides of length such that the hypotenuse is 3 inches less than twice the shorter leg and the longer leg is 3 inches longer the shorter leg. how long is each side of the triangle?
Answer by ewatrrr(24785)


Hi,
Let x represent the length of the shorter leg.
then the hypotenuse is (2x-3) and the other leg is (x+3)
Applying the Pythagorean Theorem***
x^2 + (x+3)^2 = (2x-3)^2
solving for x
x^2 + x^2 + 6x + 9 = 4x^2 -12x + 9
2x^2 - 18x = 0
2x(x-9)= 0
x = 0 Extraneious solution
x = 9, the length of the shorter leg. Other two sides are 12 & 15
CHECKING our Answer***
81 + 144 = 225